OpenStudy (anonymous):

If (70)x = 1, what are the possible values of x? Explain your answer.

OpenStudy (phi):

is that \(70^x =1 \) or \( 70 \cdot x = 1 \) ?

OpenStudy (anonymous):

Its (7^0)^x=1

OpenStudy (phi):

oh. any number with a "zero exponent" like 7^0 , is 1 in other words \[ \left(7^0\right)^x = 1^x \] 1 to any power is always 1

OpenStudy (phi):

in other words, for 1^x, x can be any number you like, and 1^x simplifies to 1

OpenStudy (anonymous):

So the value of x can be any number?

OpenStudy (phi):

yes, in your problem x can be any number, and you will always get 1

OpenStudy (anonymous):

okay, thanks